Find the f(x) = mx + b, given the x- and y-intercept of the linear function.
x-intercept -15, y-intercept -18

Answer by stanbon(75887)

Find the f(x) = mx + b, given the x- and y-intercept of the linear function.
x-intercept -15, y-intercept -18
------------------
You have 2 points:: (-15,0) and (0,-18)
----
slope = -18/(0--15) = -18/15
------
Note: y- intercept = (0,-18)
-----
Equation:
y = (-6/5)x - 18
--------------------
Cheers,
Stan H.
--------------------